I need help solving the elimination method please.
0.05x+0.25y=66
0.15x+0.05y=72

Answer by lwsshak3(11628)

I need help solving the elimination method please.
0.05x+0.25y=66
0.15x+0.05y=72
**
0.05x+0.25y=66
0.15x+0.05y=72
multiply both equations by 100
1) 5x+25y=6600
2) 15x+5y=7200
multiply eq 1 by 3
3) 15x+75y=19800
2) 15x+5y=7200
subtract eq2 from eq3
4) 70y=12600
y=12600/70=180
..
1) 5x+25y=6600
5x+25(180)=6600
5x+4500=6600
5x=2100
x=2100/5=420
..
ans:
x=420
y=180
..
Check:
.05(420)+.25(180)=21+45=66
.15(420)+.05(180)=63+9=72